Accelerator Update for August 4, 2008 - August 6, 2008
Monday August 4
The day shift began with the Tevatron (TeV) recovering from a lightning strike and resulting quench, with Operations monitoring the Recycler (RR) stash (302.3E10), with the Antiproton Source (Pbar) stacking (70.7E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) waiting for the next store, and with the SciBooNE / MiniBooNE (BNB) and NuMI experiments taking beam.
Operations reported at 8:34 AM that an RF expert worked on MIRF18 so the NuMI beam intensity could be raised back to its normal level.
At 9:28 AM, two MI kickers (I:KPS4A & B) tripped off. They reset.
At 9:33 AM, another MI kicker (KPS5S) tripped off and caused the NuMI beam permit to also trip off. Everything reset.
Operations reported at 10:08 AM the following status:
C3 Cryo System
D1 Cryo System
At 10:17 AM, a MECAR expert reported that the vertical tunes could not be read nor could changes be sent until MECAR was rebooted. However, the Recycler stash would not react well to a MECAR reboot, so this won’t be done until after the next store.
RF experts began work on MIRF2 at 10:23 AM.
At 11:11 AM, the NuMI beam permit tripped off due to MI RF. It reset.
A Linac expert reported at 11:21 AM that he was beginning to tune the driver for LRF4. Beam resumed about an hour later.
At 12:06 PM, Operations reported that two TeV quadrupole power supplies (D0Q2 and D0Q3) tripped off. A TeV expert thought the trips were due to the TeV sector D1 wet engine work.
Cryo experts reported at 12:57 PM that they had completed the D1 flywheel replacement.
Cryo system experts at CHL reported at 2:04 PM that the C3 and D1 systems were up and running.
Operations began the TeV turn on at 2:37 PM.
At 2:49 PM, KRF7 tripped off on a water flow indication. It reset.
At 3:10 PM, Operations reported losing all three MI front-end (FE) and the Tevatron Clock (TCLK). It took about 40 minutes to recover TCLK and the FEs.
MI resumed sending beam at 3:56 PM.
Operations put the TeV through a dry squeeze at 4:10 PM.
Operations reported at 4:12 PM that the Recycler had stashed antiprotons once during the day shift.
Operations put the TeV through a wet squeeze around 4:52 PM.
At 4:54 PM, Operations reported a problem with DRF1-7.
At 5:18 PM, BRF14 tripped off. It reset.
At 5:49 PM, Operations reported a rise in temperatures for the Booster 95 LCW system. Operations contacted an expert who later reported that one CUB cooling tower was down, but that probably wasn’t part of the problem.
Operations put the TeV into shot setup at 6:04 PM.
At 8:08 PM, the Booster West Anode Power Supply was found off with no indication for the trip. It reset.
At 8:10 PM, the Booster Gradient Magnet Power Supply (GMPS) was found off and would not turn on. The trip indication was a transformer overload. Operations called in experts.
At 8:22 PM, Operations reported that the Recycler had lost its ability to control Stochastic cooling. Operations called in an expert.
Operations reported at 9:16 PM that GMPS was brought back on line; Pbar resumed stacking and NuMI and BNB resumed taking beam.
Operations put the TeV through a dry squeeze at 9:40 PM.
A Recycler expert reported at 9:55 PM that he had reseated a attenuator, which returned settings control.
Operations prepared MI-31 at 9:58 PM for an access by Electron Cooling experts.
At 11:22 PM, the MI-10 BPMs stopped communicating. Operators checked the voltages, which were good, and then power cycled the BPM VME controller, with no success.
At 11:39 PM, LRF4 tripped off. It reset.
Pelletron experts reported at 11:43 PM that everything was okay, but in an odd state. Operations contacted a controls expert who suggested an ACCELNET power cycle. This restored controls and readbacks.
Operations reported at 11:58 PM that the Pelletron was cooling the beam normally.
Tuesday August 5
The midnight shift began with the TeV in a dry squeeze, with Operations monitoring the stash (352.7E10), with Pbar stacking (173.7E10), with the BNB and NuMI experiments taking beam.
At 12:03 AM, a Recycler expert reported an air conditioning problem at MI-31.
At 12:10 AM, BRF13 tripped off. It reset, but then tripped off again and would not reset. Operators bypassed the station and brought BRF19 on in its place.
At 12:39 AM, BRF19 tripped off. It reset.
Operations put the TeV into shot setup at 1:02 AM.
Operations established store 6338 at 2:56 AM with an initial luminosity of 265.77E30.
Operations reported at 3:54 AM that they had rebooted MECAR.
At 3:56 AM, the NuMI beam permit began tripping off due to MI RF problems. Nothing had tripped, but the current for MIRF17 and MIRF18 was drooping. Operations lowered the NuMI intensity.
At 5 AM, the NuMI beam permit tripped off. MIRF2 had tripped off for the first of about seven times in 14 minutes.
At 5:52 AM, a Booster power supply (B:SQL23) tripped off on a ground fault indication. It reset.
At 6:53 AM, KRF5 tripped off. It reset.
A controls expert reported at 7:19 AM that he was investigating the MI-10 BPM problems.
Operations reported at 8:05 AM that the Recycler had stashed three times during the midnight shift.
Operations reported at 8:40 AM that an RF expert was working on the MI RF. The NuMI beam intensity was returned to normal.
At 8:54 AM, TeV store 6338 aborted. There was no quench. A TeV sector C3 Dump power supply door indicated a trip. Operations asked power supply experts to investigate.
An expert reported at 9:57 AM that the C3 building knife-switch blades had scorch marks.
A RF expert reported at 10:07 AM that he was beginning to work on BRF13.
Cryo system experts reported at 10:22 AM that they were beginning work on the D3 wet engine connecting rod replacement.
An instrumentation expert reported at 10:27 AM that the MI-10 BPMs were fixed.
A TeV power supply expert began work at 10:39 AM on the F2 power supply that was taken out of the circuit over the weekend.
At 11:08 AM, the 400 MeV area tripped off for the first of two times in four minutes. It reset.
The Cryo Coordinator reported at 11:20 AM that the TeV sector D3 wet engine work was completed.
D0 came out of access at 12:27 PM.
FESS personnel reported at 12:58 PM that they had completed their work on the MI-31 AC chiller. Operators and Electron Cooling experts brought the Pelletron back online.
Operations reported at 1:22 PM that the TeV sector C3 power supply work had finished. Power supply experts put the F2 power supply back into the TeV circuit and began testing it.
Recycler experts reported at 1:23 PM that they needed to access MI-31. Operators made it ready for access.
At 2:13 PM, MIRF2 tripped off and took down the NuMI beam permit. Everything reset.
TeV power supply experts reported at 3:07 PM that they were taking the F2 power supply out of the circuit. It still had problems.
Operations put the TeV through a dry squeeze at 4:22 PM.
Operations reported at 4:27 PM that the Recycler had stashed once during the day shift.
Operations put the TeV through a wet squeeze at 4:55 PM.
Operations put the TeV through another wet squeeze at 5:43 PM.
Operations put the TeV into shot setup at 7:15 PM.
Operations established store 6341 at 9:06 PM with an initial luminosity of 270.36E30.
At 9:11 PM, MIRF2 tripped off for the first of two times in 24 minutes and took down the NuMI beam permit. Everything reset.
At 11:20 PM, KRF3 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Wednesday August 6
The midnight shift began with Operations monitoring store 6341 (155.42E30) and the stash (124.8E10), with Pbar stacking (17.1E10), and with the BNB and NuMI experiments taking beam.
At 12:10 AM, KRF3 tripped off. It reset.
At 1:31 AM, Operations reported that some glitch in Booster happened four times during the last 30 minutes that caused substantial beam position changes, which held off beam for two minutes each time it occurred.
At 5:42 AM, Operations reported a power glitch. The stack, stash, and store remained intact, but about 6E30 of luminosity were lost when TRF1 and the Tevatron Electron Lens (TEL1) tripped off. Also, all four Linac Klystron RF stations tripped off.
At 5:44 AM, all of the Booster RF stations tripped off.
At 6:25 AM, LRF4 tripped off. It reset.
At 7:14 AM, Operations reported a smoke alarm for the MI-60 service building. Operations turned off all the MI anode supplies and called in a RF expert.
Operations put Pbar into standby at 7:50 AM.
Operations reported at 7:57 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Wednesday
The plans for today are to fix MI and resume stacking, stashing, and storing.
Comments and Suggestions
|last modified 08/06/2008 email Fermilab|
|Security, Privacy, Legal|